OpenStudy (anonymous):

In the circuit shown, a voltage drop of 6.6 V wil occur across which of the following resistors? A. R_1 B. R_2 C. R_3 D. R_4

OpenStudy (anonymous):



OpenStudy (michele_laino):

here we have to compute the total or equivalent resistance

OpenStudy (michele_laino):

the equivalent resistance is given by the subsequent formula: \[\Large {R_{EQ}} = 8.5 + 3.2 + \frac{1}{{\frac{1}{{13}} + \frac{1}{{18}}}} = ...\Omega \]

OpenStudy (anonymous):

so we get 19.248?

OpenStudy (michele_laino):

ok! we can assume R_eq=19.25 ohms

OpenStudy (michele_laino):

so the total current is: I=15/19.24=...amperes

OpenStudy (anonymous):

and we get 0.779?

OpenStudy (michele_laino):

ok! I=0.779 amperes. Now the voltage drop of resistor R_1 is: V_1=8.5*0.779=...volts

OpenStudy (anonymous):

6.6215!

OpenStudy (michele_laino):

ok! maybe it is the right answer the voltage drop of resistor #4, is: V_4=3.2*0.779=...volts

OpenStudy (anonymous):

2.4928!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now the voltage drop V_AB is: 15-(6.6215+2.4928)=...volts